# A Dynamical Theory of Overfitting and Generalization in Large Two-Layer Networks

### v2

@@ Abstract
Understanding the inductive bias and generalization properties of large overparametrized machine learning models requires to characterize the dynamics of the training algorithm. We study the learning dynamics of large two-layer neural networks via dynamical mean field theory, a well established technique of non-equilibrium statistical physics. We show that, for large network width, the training dynamics exhibits a separation of timescales which implies: $(i)$ The emergence of a slow time scale associated with the growth in Gaussian/Rademacher complexity of the network; $(ii)$ Inductive bias towards small complexity if the initialization has small enough complexity; $(iii)$ A dynamical decoupling between feature learning and overfitting regimes; $(iv)$ A non-monotone behavior of the test error, associated ‘feature unlearning’ regime at large times.
@@ 1 Introduction
In contrast, in modern ML, the empirical risk is highly non-convex, the number of parameters is comparable with the number of training samples, and the model complexity is only weakly controlled. As a consequence, there are many assignments of the model parameters (many global empirical risk minimizers) that perfectly interpolate the data —even when these are noisy. While all of these interpolators have equal error on the training data, they behave very differently (and some of them very poorly) on test data. It has been hypothesized that models trained by SGD generalize well to test data because the algorithm selects a global minimum with low complexity, although a mechanistic understanding of this process is lacking. The generalization properties cannot be decoupled from the training dynamics.
@@
In this work we leverage tools from theoretical physics to directly analyze the training dynamics and derive quantitative predictions on the implicit bias of neural network training, in a simple setting. This allows us to capture feature learning and lazy/overfitting regimes within the same unified picture. We discover a time-scale separation in the training dynamics, between an early stage in which the model learns the relevant features representation of the data, and a late stage of training that is characterized by overfitting, feature ‘unlearning,’ and hence test error that increases with training. While the regularization effect of early stopping has been an important object of study (for simpler models) in the past [MB89, Bis95, ZY05, YRC07], our work is the first to point out a time-scale separation between feature learning (on a faster timescale) and overfitting (on a slower time scale), thus reconciling the feature learning and neural tangent theories of learning.
@@
where ${\boldsymbol{\theta}}=({\boldsymbol{a}},\boldsymbol{W})$, where $\boldsymbol{W}=({\boldsymbol{w}}_{1},\dots,{\boldsymbol{w}}_{m})\in{\mathbb{R}}^{d\times m}$ and ${\boldsymbol{a}}=(a_{1},\dots,a_{m})\in{\mathbb{R}}^{m}$ are, respectively, first- and second-layer weights. For convenience, we fix the normalization $\|{\boldsymbol{w}}_{i}\|=1$.
We apply model (1.1) to a supervised learning task. We are given i.i.d. data $(y_{i},{\boldsymbol{x}}_{i})$, $i\leq n$, with $y_{i}\in{\mathbb{R}}$ a response variable and ${\boldsymbol{x}}_{i}\in{\mathbb{R}}^{d}$ a feature vector, and try to learn a model $f(\,\cdot\,;{\boldsymbol{\theta}})$ to predict the response $y_{\mbox{\tiny\rm new}}$ corresponding to a new input ${\boldsymbol{x}}_{\mbox{\tiny\rm new}}$. We use gradient flow (GF) to minimize the empirical risk under square loss, namely
@@ 2 Main results: Dynamical mean field theory
Our main focus will be on the simplest case, namely $k=1$, with $\varphi$ a generic function (in particular ${\mathbb{E}}\{\varphi(G)G\}\neq 0$ for $G\sim{\sf N}(0,1)$). Some of our results apply to $k$-index models for general fixed $k$ (in particular, the rigorous results of Section 3). We defer to future work a more complete analysis of the DMFT for $k\geq 2$.
We discover a separation of time scales at large $m$ (or large $n/d$), for sufficiently small initialization $a_{0}$: feature learning takes place on a fast time scale, followed by overfitting/reversal to kernel learning. This scenario is summarized in Figure 1, which plots numerical evaluations of our theoretical predictions a $k=1$, $\tau>0$ data distribution, in the limit $n,d,m\to\infty$ at overparametrization ratio $\alpha=0.3$.
@@ 2.1 Technique
In this section we outline our results based on ‘dynamical mean field theory’ (DMFT). The next section will present rigorous results that are proven independently. Our DMFT analysis is based on the following two steps:
Step 1: We leverage techniques from theoretical physics to derive an approximate asymptotic characterization of the gradient flow dynamics (1.2) in the limit $n,d\to\infty$, with $n/d\to\overline{\alpha}$. This characterization consists of a set of integral-differential equations for the following (in probability) limits
$C_{ij}(t_{1},t_{2}):=\operatorname*{p-lim}_{n,d\to\infty}\langle{\boldsymbol{w}}_{i}(t_{1}),{\boldsymbol{w}}_{j}(t_{2})\rangle\,,\;\;\;\;{\boldsymbol{v}}_{i}(t):=\operatorname*{p-lim}_{n,d\to\infty}{\boldsymbol{U}}^{{\sf T}}{\boldsymbol{\theta}}_{i}(t)\,.$ (2.2)
@@
Here we use a DMFT that is is asymptotically exact for a well-defined Gaussian version of the original model: this allows us to solve the equations efficiently. The Gaussian approximation comes with an error which we prove is vanishing on time scales of order one and we demonstrate empirically to be small on larger time scales. The curves in Fig. 1 were obtained by solving numerically the DMFT equations, see SM.
@@
The analysis of the DMFT equations in the double limit $m,t\to\infty$ limit is an example of singular perturbation theory [Ber01, Hol13]. Making this type of analysis rigorous is notoriously challenging and we proceed by a combination of numerical solutions and analytical derivations.
In the following, we will first consider the simplest possible setting, pure noise data, and subsequently consider single-index model. The structure of the activation function and target nonlinearity will be encoded in the functions $h(q):={\mathbb{E}}\{\sigma(G_{1})\sigma(G_{q})\}$ and $\widehat{\varphi}(q):={\mathbb{E}}\{\varphi(G_{1})\sigma(G_{q})\}$, where $G_{1},G_{q}$ are standard jointly Gaussian with ${\mathbb{E}}\{G_{1}G_{q}\}=q$.
As mentioned above, we assume throughout $n,d\to\infty$, with $n/d\to\overline{\alpha}\in(0,\infty)$, with the limit $m,\overline{\alpha}\to\infty$ taken afterwards. To further simplify our analysis, we assume a symmetric initialization whereby $a_{i}(0)=a_{0}$ is independent of $i\leq m$ and $({\boldsymbol{w}}_{i}(0):i\leq m)\sim_{iid}{\sf Unif}({\mathbb{S}}^{d-1})$.
@@ 2.2 Training on pure noise
Figure 2: Evolution of second-layer weights (left) and train error (right) when fitting pure noise data. Here we use mean field initialization and $h(z)=(9/10)z+(1/6)z^{3}$ and $\tau=0.6$. Symbols: SGD results on actual 2-layer networks. Continuous viridis lines: Numerical solution of the DMFT equations.
@@
The ability of a model ${\mathcal{F}}_{\Theta}=(f(\,\cdot\,;{\boldsymbol{\theta}}):{\boldsymbol{\theta}}\in\Theta)$ to interpolate pure noise is intimately connected to its Gaussian complexity ${\mathcal{G}}({\mathcal{F}}_{\Theta};n):={\mathbb{E}}\sup_{{\boldsymbol{\theta}}\in\Theta}\langle{\boldsymbol{g}},f(\boldsymbol{X};{\boldsymbol{\theta}})\rangle/n$ [Ver18]. Indeed, interpolation is impossible unless ${\mathcal{G}}({\mathcal{F}}_{\Theta};n)\geq\tau$. Viceversa, ${\mathcal{G}}({\mathcal{F}}_{\Theta};n)\ll\tau$ ensures good generalization.
@@
In the SM, we will make precise the above picture of the evolution of $a(t)$. Here, we consider a simpler setting in which $a(t)=\gamma\sqrt{m}$ with $\gamma$ independent of $m$, not evolving with training. Note that ${\mathcal{G}}({\mathcal{F}}_{\Theta};n)\asymp\gamma/\sqrt{\alpha}$ and hence such a network can interpolate pure noise if $\gamma$ is larger than threshold depending on $\alpha$. Our DMFT predicts a phase transition. For $\alpha\in(0,1)$, GF converges to vanishing train error with high probability if $\gamma>\gamma_{\mbox{\tiny\rm GF}}(\alpha,m)\tau$, and converges to a strictly positive training error if $\gamma<\gamma_{\mbox{\tiny\rm GF}}(\alpha,m)\tau$. The threshold $\gamma_{\mbox{\tiny\rm GF}}(\alpha,m)$ converges to a limit $\gamma_{\mbox{\tiny\rm GF}}^{*}(\alpha)\in(0,1)$ as $m\to\infty$.
@@
Informally $\gamma_{\mbox{\tiny\rm GF}}^{*}(\alpha)$ is the minimum complexity $\gamma$ for a very large network to interpolate noise. The functions $\gamma_{\mbox{\tiny\rm GF}}^{*}(\alpha)$, $e_{*}(\gamma)$ will play an important role below.
We next consider training on data rom a single-index model. The initial scale of second-layer weights $\|{\boldsymbol{a}}(0)/m\|_{1}$ plays a crucial role and we will separately analyze lazy and mean field initializations.
@@ 2.3 Training on data with latent structure: lazy initialization
Figure 3: Train/test error (right) when fitting data from a single index model. We set $h(z)=\widehat{\varphi}(z)=(9/10)z+z^{2}/2$, $\tau=0.3$ and $\alpha=0.3$. Lines correspond to predictions from the DMFT (continuous: train error; dashed: test error). The inset on the right reports the value of the test error at the plateau, as a function of $m$ (green line is the $m\to\infty$ value). Right: Same data plotted versus $t$.
@@
$\lim_{n,d\to\infty}\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{\theta}}(t))=e_{\mbox{{\rm\footnotesize tr}}}(t;\varphi,\gamma_{0},m,\alpha)\,,\;\;\;\;\;\lim_{n,d\to\infty}\mathscrsfs{R}({\boldsymbol{\theta}}(t))=e_{\mbox{\rm\footnotesize ts}}(t;\varphi,\gamma_{0},m,\alpha)\,.$ (2.4)
@@
For large scaled time $\hat{t}$, the error ${e}^{\mbox{\tiny\rm lz1}}(\hat{t};\varphi,\gamma_{0},\alpha)$ converges to the error of the best linear approximation to $f_{*}$. This dynamical regime follows the qualitative predictions of NTK theory, and is essentially linear in the weights ${\boldsymbol{w}}_{i}$.
Second dynamical regime: $t=\Theta(1)$. Second layer weights do not change significantly: $\gamma(t)=\gamma_{0}+o_{m}(1)$, while first layer weights change significantly $\|{\boldsymbol{w}}_{i}(t)-{\boldsymbol{w}}_{i}(0)\|=\Theta(1)$. However they change orthogonally to the latent subspace $\boldsymbol{U}$} and hence the test error does not change: no actual learning takes place in this regime.
@@
Third dynamical regime: $t=\Theta(m)$. The qualitative properties of this regime depend whether or not $\gamma_{0}$ is larger than an interpolation threshold $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, which generalizes the threshold $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)=\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,0,1)$ introduced in the pure noise case. Because of the equivalence between the two dynamics mentioned above, the following relation exists:
@@
For $\gamma_{0}>\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, interpolation is achieved during the second dynamical regime, no further evolution takes place. For $\gamma_{0}<\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, we have, for any $z\in(0,\infty)$, as $m\to\infty$,
$\gamma(mz)={\gamma}^{\mbox{\tiny\rm lz3}}(z)+o_{m}(1),\;\;\;e_{\mbox{{\rm\footnotesize tr}}}(mz)={e}^{\mbox{\tiny\rm lz3}}_{\mbox{{\rm\footnotesize tr}}}(z)+o_{m}(1),\;\;\;e_{\mbox{\rm\footnotesize ts}}(mz)={e}^{\mbox{\tiny\rm lz3}}_{\mbox{\rm\footnotesize ts}}(z)+o_{m}(1)\,$, (2.8)
with ${\gamma}^{\mbox{\tiny\rm lz3}}(z)$ growing to $\overline{\gamma}^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)\approx\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$ and ${e}^{\mbox{\tiny\rm lz3}}_{\mbox{{\rm\footnotesize tr}}}(z)$ decreasing to 0 as $z\to\infty$. In other words, interpolation is achieved on this third regime. Further ${e}^{\mbox{\tiny\rm lz3}}_{\mbox{\rm\footnotesize ts}}(z)$ increase from $\gamma_{0}$ approximately to the test error predicted in Eq. (2.6), with $\gamma_{0}$ replaced by $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$.

### v3

# Dynamical Decoupling of Generalization and Overfitting in Large Two-Layer Networks

@@ Abstract
Understanding the inductive bias and generalization properties of large overparametrized machine learning models requires to characterize the dynamics of the training algorithm. We study the learning dynamics of large two-layer neural networks via dynamical mean field theory, a well established technique of non-equilibrium statistical physics. We show that, for large network width $m$, and large number of samples per input dimension $n/d$, the training dynamics exhibits a separation of timescales which implies: $(i)$ The emergence of a slow time scale associated with the growth in Gaussian/Rademacher complexity of the network; $(ii)$ Inductive bias towards small complexity if the initialization has small enough complexity; $(iii)$ A dynamical decoupling between feature learning and overfitting regimes; $(iv)$ A non-monotone behavior of the test error, associated ‘feature unlearning’ regime at large times.
@@ 1 Introduction
In contrast, in modern ML, the empirical risk is highly non-convex, the number of parameters is comparable with the number of training samples, and the model complexity is only weakly controlled. As a consequence, there can be many assignments of the model parameters (many global empirical risk minimizers) that perfectly interpolate the data —even when these are noisy. While all of these interpolators are indistinguishable on the training data, they behave very differently (and some of them very poorly) on test data. It has been hypothesized that models trained by SGD generalize well to test data because the algorithm selects a near global minimizer with low complexity, although a mechanistic understanding of this process is lacking. For this reason, the generalization properties cannot be decoupled from the training dynamics.
@@
In this work we leverage tools from theoretical physics to directly analyze the training dynamics and derive quantitative predictions on the implicit bias of neural network training, in a simple setting. This allows us to capture feature learning and lazy/overfitting regimes within the same unified picture. We discover a time-scale separation in the training dynamics, between an early stage in which the model learns the relevant features representation of the data, and a late stage of training that is characterized by overfitting, feature ‘unlearning,’ and hence test error that increases with training. While the regularizing effect of early stopping has been an important object of study (for simpler models) in the past [MB89, Bis95, ZY05, YRC07], our work is the first to point out a time-scale separation between feature learning (on a faster timescale) and overfitting (on a slower time scale), thus reconciling the feature learning and neural tangent theories of learning.
@@
where ${\boldsymbol{\theta}}=({\boldsymbol{a}},\boldsymbol{W})$, where $\boldsymbol{W}=({\boldsymbol{w}}_{1},\dots,{\boldsymbol{w}}_{m})\in{\mathbb{R}}^{d\times m}$ and ${\boldsymbol{a}}=(a_{1},\dots,a_{m})\in{\mathbb{R}}^{m}$ are, respectively, first- and second-layer weights. For convenience, we fix the normalization $\|{\boldsymbol{w}}_{i}\|=1$, and assume that $\sigma$ does not depend on $m$. We apply model (1.1) to a supervised learning task. We are given i.i.d. data $(y_{i},{\boldsymbol{x}}_{i})$, $i\leq n$, with $y_{i}\in{\mathbb{R}}$ a response variable and ${\boldsymbol{x}}_{i}\in{\mathbb{R}}^{d}$ a feature vector, and try to learn a model $f(\,\cdot\,;{\boldsymbol{\theta}})$ to predict the response $y_{\mbox{\tiny\rm new}}$ corresponding to a new input ${\boldsymbol{x}}_{\mbox{\tiny\rm new}}$. We use gradient flow (GF) to minimize the empirical risk under square loss, namely
@@ 2 Main results: Dynamical mean field theory
Our main focus will be on the simplest case, namely $k=1$, with $\varphi$ a generic function (in particular ${\mathbb{E}}\{\varphi(G)G\}\neq 0$ for $G\sim{\sf N}(0,1)$, which corresponds information exponent equal to one according to the classification of [AGJ21].). Some of our results apply to $k$-index models for general fixed $k$ (in particular, the rigorous results of Section 3). We defer to future work a more complete analysis of the DMFT for $k\geq 2$.
We discover a separation of time scales at large $m$ (or large $n/d$), for sufficiently small initialization $a_{0}$: feature learning takes place on a fast time scale, followed by overfitting/reversal to kernel learning. This scenario is summarized in Figure 1, which plots numerical evaluations of our theoretical predictions at $k=1$, $\tau>0$ data distribution, in the limit $n,d,m\to\infty$ at overparametrization ratio $\alpha=0.3$.
@@
In this section we outline our results based on ‘dynamical mean field theory’ (DMFT). The next section will present rigorous results that are proven independently.
@@ 2.1 Technique
Our DMFT analysis is based on the following two steps:
Step 1: We leverage techniques from theoretical physics to derive an approximate asymptotic characterization of the gradient flow dynamics (1.2) in the limit $n,d\to\infty$, with $n/d\to\overline{\alpha}$. This characterization consists of a set of integral-differential equations for the following asymptotic quantities (here $\operatorname*{p-lim}$ denotes limit in probability, and we use the superscripts $n$ to emphasize the dependence of the right-hand side on $n,d$)
$\begin{split}&C_{ij}(t_{1},t_{2}):=\operatorname*{p-lim}_{n,d\to\infty}\langle{\boldsymbol{w}}^{n}_{i}(t_{1}),{\boldsymbol{w}}^{n}_{j}(t_{2})\rangle\,,\\ &{\boldsymbol{v}}_{i}(t):=\operatorname*{p-lim}_{n,d\to\infty}{\boldsymbol{U}}^{{\sf T}}{\boldsymbol{w}}^{n}_{i}(t)\,,\;\;\;\;a_{i}(t):=\operatorname*{p-lim}_{n,d\to\infty}a^{n}_{i}(t)\,.\end{split}$ (2.2)
@@
However, the asymptotically exact DMFT characterization of [CCM21] is rather complex to integrate numerically or to study analytically. In order to circumvent this problem, we use a DMFT that is is asymptotically exact for a well-defined Gaussian version of the original model. Namely, we observe that the empirical risk of Eq. (1.2) takes the form
$\widehat{\mathscrsfs{R}}_{n}({\boldsymbol{\theta}})=\frac{1}{2n}\big\|{\boldsymbol{F}}({\boldsymbol{\theta}})\big\|^{2}\,$, (2.3)
where ${\boldsymbol{F}}:({\mathbb{S}}^{d-1})^{m}\times{\mathbb{R}}^{m}\to{\mathbb{R}}^{n}$ is s stochastic process with i.i.d. components $F_{i}({\boldsymbol{\theta}})=y_{i}-f({\boldsymbol{x}}_{i};{\boldsymbol{\theta}})$. We replace these by Gaussian processes with matching mean and covariance, and study the DMFT for gradient flow with respect to the associated risk $\widehat{\mathscrsfs{R}}^{g}_{n}({\boldsymbol{\theta}})$.
The Gaussian approximation comes with an error which we show analytically is vanishing on time scales of order one ( indeed on these time scales we correctly recover the mean field theory of [MMN18, CB18]) and we demonstrate empirically to be small on larger time scales ( see for instance example Fig. 4.) The curves in Fig. 1 were obtained by solving numerically the DMFT equations, see Appendix C for details.
@@
The analysis of the DMFT equations in the double limit $m,t\to\infty$ is an example of singular perturbation theory [Ber01, Hol13]. Making this type of analysis rigorous is notoriously challenging and we proceed by a combination of numerical solutions and analytical derivations.
In the following, we will first consider the simplest possible setting, pure noise data, and subsequently consider the single-index model. The structure of the activation function and target nonlinearity will be encoded in the functions
$h(q):={\mathbb{E}}\{\sigma(G_{1})\sigma(G_{q})\},\;\;\;\;\widehat{\varphi}(q):={\mathbb{E}}\{\varphi(G_{1})\sigma(G_{q})\}\,$,
where $G_{1},G_{q}$ are standard jointly Gaussian with ${\mathbb{E}}\{G_{1}G_{q}\}=q$. The relation between $\sigma,\varphi$ and $h,\widehat{\varphi}$ is conveniently expressed in terms of the expansions in Hermite polynomials $\sigma(x)=\sum_{k\geq 0}s_{k}{\rm He}_{k}(x)$, $\varphi(x)=\sum_{k\geq 0}f_{k}{\rm He}_{k}(x)$, which corresponds to the analytic expansion $h(q)=\sum_{k\geq 0}s_{k}^{2}q^{k}$, $\widehat{\varphi}(q)=\sum_{k\geq 0}s_{k}f_{k}q^{k}$.
As mentioned above, we assume throughout $n,d\to\infty$, with $n/d\to\overline{\alpha}\in(0,\infty)$, with the limit $m,\overline{\alpha}\to\infty$ taken afterwards. To further simplify our analysis, we assume a symmetric initialization whereby $a_{i}(0)=a_{0}$ is independent of $i\leq m$ and $({\boldsymbol{w}}_{i}(0):i\leq m)\sim_{iid}{\sf Unif}({\mathbb{S}}^{d-1})$. Throughout, we use ‘with high probability’ for ‘with probability converging to one as $n,d\to\infty$.’
@@ 2.2 Training on pure noise
Figure 2: Evolution of second-layer weights (left) and train error (right) when fitting pure noise data. Here we use mean field initialization, $h(z)=(9/10)z+(1/6)z^{3}$, $\alpha=0.4$ and $\tau=0.6$. Symbols: SGD results on actual 2-layer networks with $d=200$, $n=\alpha md$ (averaged over 10 simulations). Continuous viridis lines: Numerical solution of the DMFT equations. Note that the second layer weights are given in terms of a scalar quantity as the result of the statistically symmetric initialization.
@@
The ability of a model ${\mathcal{F}}_{\Theta}=(f(\,\cdot\,;{\boldsymbol{\theta}}):{\boldsymbol{\theta}}\in\Theta)$ to interpolate pure noise is intimately connected to its Gaussian complexity ${\mathcal{G}}({\mathcal{F}}_{\Theta};n):={\mathbb{E}}\sup_{{\boldsymbol{\theta}}\in\Theta}\langle{\boldsymbol{g}},f(\boldsymbol{X};{\boldsymbol{\theta}})\rangle/n$ [Ver18] (where ${\boldsymbol{g}}\sim{\sf N}({\mathbf{0}},{\boldsymbol{I}}_{n})$ is independent of $f(\boldsymbol{X},;{\boldsymbol{\theta}})=(f({\boldsymbol{x}}_{i};{\boldsymbol{\theta}}):\,i\leq n)$. Indeed, interpolation is impossible unless ${\mathcal{G}}({\mathcal{F}}_{\Theta};n)\geq\tau$. Viceversa, ${\mathcal{G}}({\mathcal{F}}_{\Theta};n)\ll\tau$ ensures good generalization.
@@
In Section G.1 of the appendix, we will make precise the above picture of the evolution of $a(t)$. Here, we consider a simplified setting in which $a(t)=\gamma\sqrt{m}$ with $\gamma$ independent of $m$, not evolving with training. Note that ${\mathcal{G}}({\mathcal{F}}_{\Theta};n)\asymp\gamma/\sqrt{\alpha}$ and hence such a network can interpolate pure noise if $\gamma$ is larger than threshold depending on $\alpha$. Our DMFT predicts a sharp phase transition. For $\alpha\in(0,1)$, GF converges to vanishing train error with high probability if $\gamma>\gamma_{\mbox{\tiny\rm GF}}(\alpha,m)\tau$, and converges to a strictly positive training error if $\gamma<\gamma_{\mbox{\tiny\rm GF}}(\alpha,m)\tau$. The threshold $\gamma_{\mbox{\tiny\rm GF}}(\alpha,m)$ converges to a limit $\gamma_{\mbox{\tiny\rm GF}}^{*}(\alpha)\in(0,1)$ as $m\to\infty$.
@@
Informally $\gamma_{\mbox{\tiny\rm GF}}^{*}(\alpha)$ is the minimum complexity $\gamma$ for a very large network to interpolate noise via gradient flow. The functions $\gamma_{\mbox{\tiny\rm GF}}^{*}(\alpha)$, $e_{*}(\gamma)$ will play an important role below.
We will next consider training on data from a single-index model. The initial scale of second-layer weights $\|{\boldsymbol{a}}(0)/m\|_{1}$ plays a crucial role and we will separately analyze lazy and mean field initializations.
@@ 2.3 Training on data with latent structure: lazy initialization
Figure 3: Train/test error (right) when fitting data from a single index model. We set $h(z)=\widehat{\varphi}(z)=(9/10)z+z^{2}/2$, $\tau=0.3$ and $\alpha=0.3$. Lines correspond to predictions from the DMFT (continuous: train error; dashed: test error). Black continuous line is the $m\to\infty$ value. Right: Same data plotted versus $t$.
@@
$\lim_{n,d\to\infty}\widehat{\mathscrsfs{R}}_{n}^{g}({\boldsymbol{\theta}}(t))=e_{\mbox{{\rm\footnotesize tr}}}(t;\varphi,\gamma_{0},m,\alpha)\,,\;\;\;\;\;\lim_{n,d\to\infty}\mathscrsfs{R}^{g}({\boldsymbol{\theta}}(t))=e_{\mbox{\rm\footnotesize ts}}(t;\varphi,\gamma_{0},m,\alpha)\,.$ (2.5)
@@
For large scaled time $\hat{t}$, the error ${e}^{\mbox{\tiny\rm lz1}}(\hat{t};\varphi,\gamma_{0},\alpha)$ converges to the error of the best linear approximation to $f_{*}$. This dynamical regime follows the qualitative predictions of NTK theory, and is essentially linear in the weights ${\boldsymbol{w}}_{i}$, but the time is too short for the model to overfit the data.
Second dynamical regime: $t=\Theta(1)$. Second layer weights do not change significantly: $\gamma(t)=\gamma_{0}+o_{m}(1)$, while first layer weights change significantly $\|{\boldsymbol{w}}_{i}(t)-{\boldsymbol{w}}_{i}(0)\|=\Theta(1)$. However they change orthogonally to the latent subspace $\boldsymbol{U}$} and hence the test error does not change: no actual learning takes place in this regime, but the model starts to overfit the data.
@@
Third dynamical regime: $t=\Theta(m)$. The qualitative properties of this regime depend whether or not $\gamma_{0}$ is larger than an interpolation threshold $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, which generalizes the threshold $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)=\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,0,1)$ introduced in the pure noise case. Because the dynamics of weights ${\boldsymbol{w}}_{i}$ in the subspace orthogonal to $\boldsymbol{U}$} is equivalent to dynamics in pure noise, we expect the interpolation threshold $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$ to be given in terms of pure noise threshold $\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha)$ as follows:
@@
For $\gamma_{0}>\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, interpolation is achieved during the second dynamical regime, no further evolution takes place.
For $\gamma_{0}<\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, a non-trivial evolution takes place for $t=\Theta(m)$. Introducing the rescaled time $z\in(0,\infty)$, we obtain, as $m\to\infty$,
$\gamma(mz)={\gamma}^{\mbox{\tiny\rm lz3}}(z)+o_{m}(1),\;\;\;e_{\mbox{{\rm\footnotesize tr}}}(mz)={e}^{\mbox{\tiny\rm lz3}}_{\mbox{{\rm\footnotesize tr}}}(z)+o_{m}(1),\;\;\;e_{\mbox{\rm\footnotesize ts}}(mz)={e}^{\mbox{\tiny\rm lz3}}_{\mbox{\rm\footnotesize ts}}(z)+o_{m}(1)\,.$ (2.9)
Further, for large values of the rescaled time $z\to\infty$, ${\gamma}^{\mbox{\tiny\rm lz3}}(z)$ grows to $\overline{\gamma}^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)\approx\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$, while ${e}^{\mbox{\tiny\rm lz3}}_{\mbox{{\rm\footnotesize tr}}}(z)$ decreases to 0. In other words, interpolation is achieved on this third regime.
Further the test error ${e}^{\mbox{\tiny\rm lz3}}_{\mbox{\rm\footnotesize ts}}(z)$ increases from ${e}^{\mbox{\tiny\rm lz2}}_{\mbox{\rm\footnotesize ts}}(t;\varphi,\gamma_{0},\alpha)$ to ${e}^{\mbox{\tiny\rm lz2}}_{\mbox{\rm\footnotesize ts}}(t;\varphi,\gamma^{*}_{\mbox{\tiny\rm GF}},\alpha)$, with $\gamma^{*}_{\mbox{\tiny\rm GF}}=\gamma^{*}_{\mbox{\tiny\rm GF}}(\alpha,\varphi,\tau)$ whereby ${e}^{\mbox{\tiny\rm lz2}}_{\mbox{\rm\footnotesize ts}}(\cdots)$ is given by Eq. (2.7).